### v3

@@ 5.3 Crop Yield Prediction Experiments
For crop yield prediction, we selected the U.S. Corn Belt as the study region, the world’s largest corn-producing area with rich historical yield data. The model training incorporated three data types: remote sensing products, meteorological observations, and crop yield records. All data were sourced from the USDA National Agricultural Statistics Service (NASS, 2020).
The experiment comprised both local (intra-region) and cross-domain (inter-region) prediction tasks. We compared four models: Random Forest (RF), Deep Neural Network (DNN), Domain-Adversarial Neural Network (DANN), and an improved version (ADANN).
Figure 16: Corn planting counties in two ecosystem regions (Eastern Temperate Forest (ETF) and Great Plains (GP)) used in cross-domain prediction experiments.
@@
In addition, inter-annual environmental variability significantly affected model performance. For instance, abnormal wet conditions in parts of the Corn Belt in 2019 weakened the correlation between remote sensing features and actual yields, exacerbating domain discrepancies and degrading traditional model performance. However, ADANN maintained stable and accurate predictions under these extreme conditions, showcasing its strong robustness and adaptability to real-world variations.
Overall, the four-year experimental results indicate that ADANN consistently performs well across both local and cross-domain settings. In particular, its superior accuracy and lower error margins in domain transfer tasks underscore its suitability for cross-region yield modeling. Traditional models, while effective in localized scenarios, exhibit significant performance degradation in cross-domain applications due to their inability to adapt to domain shifts. ADANN, by contrast, emerges as the most effective solution for cross-domain agricultural remote sensing yield prediction, thanks to its stability, generalization ability, and adaptability.
@@
In summary, ADANN effectively extracts domain -invariant features through adversarial training, mitigating distribution shifts and reducing prediction bias. Its superior performance across four consecutive years in the U.S. Corn Belt validates its practicality and robustness. By maintaining a favorable balance between adversarial domain alignment and prediction error minimization, ADANN delivers stable and accurate cross-domain yield predictions, setting a new benchmark for agricultural remote sensing modeling.
@@ 5.4 Remote Sensing-Based Agricultural Land Extraction Experiments
In the comparative experiment for agricultural land extraction from remote sensing images, we selected six representative domain adaptation methods for performance evaluation, including the baseline model (trained solely on the source domain without adaptation), AdaptSegNet [101], ADVENT [102], BDL [103], IntraDA [66], and TransFusion-DualDA [85]. The evaluation was conducted across three benchmark agricultural datasets that vary significantly in spatial resolution and scene complexity:
  • DeepGlobe: High-resolution imagery at 0.5 meters per pixel, containing fine-grained farmland boundary annotations;
  • LoveDA: Ultra-high-resolution (0.3 m/pixel) scenes from urban-rural transition zones with heterogeneous land use;
  • GID: Moderate-resolution (4 m/pixel) dataset covering large regional farmland distributions.
@@
With ADVENT (d) and IntraDA (e), visual quality improves. ADVENT leverages adversarial learning to mitigate domain gaps, while IntraDA introduces intra-class consistency constraints to improve boundary smoothness. However, both methods still exhibit issues such as boundary blurring and fragmented regions.
TransFusion-DualDA (f) produces the most consistent and precise results, with boundaries closely matching ground truth and accurate region-level predictions. This performance is attributed to its long-range dependency modeling and all-scale feature fusion, which enable simultaneous capture of fine details and global structure. Nonetheless, minor edge artifacts and noise remain, suggesting room for further enhancement in handling high-resolution image heterogeneity.
In conclusion, TransFusion-DualDA demonstrates superior generalization across different spatial resolutions and land types. Its advantages in COM and IoU metrics confirm its robustness in shape-preserving segmentation tasks, making it a powerful tool for unsupervised domain adaptation in remote sensing-based agricultural land monitoring.
@@ 5.5 Summary of Experimental Findings
In the plant disease transfer detection task, MSUN demonstrates a leading performance with an average accuracy of 56.06%, surpassing other approaches such as DSAN (42.87%). It particularly excels in the cross-domain classification of corn, grape, sweet pepper, potato, and apple diseases, indicating strong adaptability to complex agricultural scenarios. Although its detection accuracy for tomato diseases is slightly lower (48.67%), models like DANN (49.51%) and MRAN (49.34%) offer potential optimization baselines. Notably, sweet pepper classification yields high accuracy across all methods (above 84%), due to its distinct morphological features.
The superior performance of MSUN stems from its innovative architectural design. Multi-representation modules are employed to extract diverse features, reducing both domain and category discrepancies. Sub-domain adaptation enhances fine-grained semantic alignment, while uncertainty regularization effectively suppresses the noise introduced by pseudo-labels. These integrated mechanisms enable MSUN to achieve optimal results on four public benchmark datasets, positioning it as a preferred solution for multi-crop disease diagnosis with improved generalization and accuracy.
@@
Model interpretability analysis suggests that ADANN effectively balances domain loss and prediction loss to mitigate the effects of domain shift. Error spatial distribution analysis (Figure 17) highlights that RF and DNN exhibited strong error clustering in regions like Michigan and Ohio. While DANN partially alleviates this through adversarial training, ADANN further optimizes prediction errors in core agricultural zones such as Illinois, underscoring its robustness under dynamic cross-region shifts.
@@
Overall, the experiments validate the effectiveness of domain-adversarial training and entropy-guided sub-domain partitioning in TransFusion-DualDA, especially under cross-sensor and cross-region conditions. This method significantly reduces the dependency on manual annotations by narrowing the domain gap in complex environments such as urban-rural transitions. Future work will focus on optimizing global feature extraction in high-resolution scenes to further enhance adaptability in dynamic remote sensing applications.
@@ 6 Challenges and Future Research Directions
Current research in agricultural intelligence confronts multifaceted challenges, spanning data complexity, methodological adaptability, and technological limitations. While domain adaptation (DA) techniques have shown promise in alleviating cross-domain issues in agricultural image analysis—such as environmental heterogeneity and annotation scarcity—they still face significant bottlenecks in computational efficiency, model interpretability, and generalizability under complex field scenarios. This section highlights key practical pain points and proposes targeted research directions.
@@ 6.1 Model Adaptability Barriers in Field Scenarios
The application of deep models in real-world agricultural environments is hindered by the following core issues:
1) Multi-scale spatiotemporal data fusion: There are significant resolution disparities between satellite imagery (10–100 m/pixel), drone imagery (0.1–1 m/pixel), and ground sensor data (centimeter-level). Existing 3D convolutional approaches struggle to achieve effective cross-scale feature alignment.
2) Severe labeled data scarcity: Field datasets often contain fewer than 50 samples per class, especially for rare crop diseases. Traditional transfer learning models exhibit a performance drop exceeding 30% when training sample sizes fall below 100.
3) Limitations in multimodal collaboration: Hyperspectral imaging in field settings suffers from radiometric fluctuations (±30% in light intensity), which induce distributional shifts across time periods. Such variance reduces the effectiveness of direct multimodal integration.
@@ 6.2 Key Future Research Directions
To overcome these challenges, four critical research directions are proposed:
(1) Lightweight Spatiotemporal Modeling Framework: Develop ultra-compact network architectures with fewer than 1M parameters, incorporating physical priors (e.g., crop growth models) to guide feature extraction. Emphasize spatiotemporal decoupling to enable cross-scale fusion between drone (5 cm) and satellite (10 m) imagery. Aim to reduce computational energy consumption to 20% of current 3D CNN baselines.
(2) Weakly Supervised Transfer Learning Paradigm: Construct pseudo-labeling mechanisms leveraging temporal continuity in crop growth stages to generate self-supervised learning signals. Implement curriculum learning strategies informed by phenological patterns to maintain over 85% classification accuracy with under 100 labeled samples.
(3) Robust Multimodal Fusion System: Design spectral normalization-based feature extraction to suppress illumination-related noise. Develop physically informed adversarial training to ensure cross-modality consistency under ±40% lighting changes. Address cross-view spatial alignment between aerial and ground sensors by establishing sub-meter correspondence standards.
(4) Privacy-Preserving Collaborative Learning: Enhance communication efficiency in federated learning frameworks to support daily updates within 10 Mbps bandwidth. Introduce differential privacy strategies compliant with GDPR, maintaining performance degradation below 5% with a privacy budget of $\epsilon=2$.
To support real-world deployment, a standardized open benchmark is recommended, including over 10 typical crops across three or more climate zones. The benchmark should incorporate field-specific noise factors (e.g., device error ±5%, occlusion rate 20%) to ensure alignment between algorithmic development and application robustness.
@@ 7 Conclusion
Adaptive modeling in agriculture faces persistent challenges, particularly in handling the complexity of spatiotemporal dynamics, the scarcity of annotated samples, and the heterogeneity of multimodal data sources. As agricultural environments grow more complex, the effectiveness of existing domain adaptation techniques in field scenarios—especially those involving cross-seasonal, cross-regional, and cross-modal variability—remains limited.
Future research should focus on building task-specific high-dimensional adaptive models to enhance performance in diverse agricultural applications, such as crop health monitoring, yield forecasting, and pest recognition. Unsupervised domain adaptation is especially valuable under annotation-scarce conditions, while multimodal fusion —including image, weather, and soil data—holds potential for further improving model generalization. Moreover, source-free domain adaptation represents a critical frontier where access to source data is restricted due to privacy or operational constraints.
The field-deployed robustness of these models hinges on their ability to operate under unpredictable environmental conditions, diverse crop types, and varying sensor platforms. Addressing these technical and operational barriers is essential for advancing precision agriculture and intelligent farming systems. Improved cross-domain adaptability and model resilience will ultimately support more accurate agricultural decision-making, yield estimation, and disease surveillance, contributing to the transformation of modern agricultural practices.

### v4

@@ 5.4 Crop Yield Prediction Experiments
For crop yield prediction, we selected the U.S. Corn Belt as the study region—the world’s largest corn-producing area with abundant historical yield data. Model training incorporated three types of data: remote sensing products, meteorological observations, and crop yield records. All data were sourced from the USDA National Agricultural Statistics Service (NASS, 2020), as shown in Figure 16. The experiment included both local (intra-region) and cross-domain (inter-region) prediction tasks. We compared four models: Random Forest (RF), Deep Neural Network (DNN), Domain-Adversarial Neural Network (DANN), and an improved variant (ADANN).
Figure 16: Corn planting counties in two ecosystem regions—Eastern Temperate Forest (ETF) and Great Plains (GP)—used in cross-domain prediction experiments.Image adapted from Ma et al.  [141].
@@
Furthermore, inter-annual environmental variability significantly affected model performance. For instance, abnormal wet conditions in parts of the Corn Belt in 2019 weakened the correlation between remote sensing features and actual yields, exacerbating domain discrepancies and degrading the performance of traditional models. Nonetheless, ADANN maintained stable and accurate predictions under such extreme conditions, demonstrating robust adaptability to real-world variations.
Overall, the four-year experimental results indicate that ADANN consistently performs well across both local and cross-domain settings. Its superior accuracy and lower error margins in domain transfer tasks underscore its suitability for cross-region yield modeling. While traditional models are effective in localized scenarios, they suffer substantial performance degradation under domain shifts. By contrast, ADANN emerges as a reliable solution for cross-domain agricultural remote sensing yield prediction, due to its stability, generalization capability, and adaptability.
@@
In summary, ADANN effectively extracts domain-invariant features through adversarial training, mitigating distribution discrepancies and reducing prediction bias. Its superior performance across four consecutive years in the U.S. Corn Belt validates its practicality and robustness. By maintaining a favorable balance between adversarial domain alignment and prediction accuracy, ADANN delivers stable and precise cross-domain yield predictions, setting a new benchmark for agricultural remote sensing modeling.
@@ 5.5 Remote Sensing-Based Agricultural Land Extraction Experiments
In the comparative experiment on agricultural land extraction from remote sensing imagery, we selected six representative domain adaptation methods for performance evaluation: a baseline model (trained solely on the source domain without adaptation), AdaptSegNet [102], ADVENT [103], BDL [104], IntraDA [67], and TransFusion-DualDA [86], as shown in Figure 18. The evaluation was conducted on three benchmark datasets (DeepGlobe, LoveDA, and GID), which differ significantly in spatial resolution and scene complexity.
@@
With ADVENT (d) and IntraDA (e), segmentation quality improves. ADVENT employs adversarial learning to reduce domain gaps, while IntraDA introduces intra-class consistency constraints to enhance boundary smoothness. However, both methods still suffer from issues such as boundary blurring and region fragmentation.
TransFusion-DualDA (f) produces the most consistent and accurate segmentation results, with boundaries closely aligning to the ground truth and precise region-level predictions. Its advantage arises from effective long-range dependency modeling and all-scale feature fusion, enabling the simultaneous capture of fine details and global structure. Nevertheless, minor edge artifacts and residual noise remain, indicating opportunities for further enhancement in managing high-resolution image heterogeneity.
In conclusion, TransFusion-DualDA exhibits superior generalization across diverse spatial resolutions and land types. Its consistent advantages in COM and IoU metrics affirm its robustness for shape-preserving segmentation, making it a highly effective method for unsupervised domain adaptation in remote sensing-based agricultural land monitoring.
@@ 5.6 Summary of Experimental Findings
In the plant disease transfer detection task, MSUN achieves a leading average accuracy of 56.06%, substantially outperforming alternative methods such as DSAN (42.87%). It particularly excels in cross-domain classification tasks involving corn, grape, sweet pepper, potato, and apple diseases, demonstrating strong adaptability to complex agricultural scenarios. Although MSUN’s accuracy on tomato disease classification is slightly lower (48.67%), models like DANN (49.51%) and MRAN (49.34%) provide potential baselines for further optimization. Notably, sweet pepper classification yields high accuracy across all methods (above 84%), likely due to its distinct morphological features.
The superior performance of MSUN can be attributed to its innovative architectural design. Multi-representation modules are employed to extract diverse features, thereby reducing both domain and category discrepancies. Sub-domain adaptation facilitates fine-grained semantic alignment, while uncertainty regularization effectively suppresses the noise introduced by pseudo-labels. These integrated mechanisms enable MSUN to achieve optimal results on four public benchmark datasets, positioning it as a preferred solution for multi-crop disease diagnosis with enhanced generalization and accuracy.
@@
Interpretability analysis reveals that ADANN effectively balances domain and prediction losses to mitigate the effects of domain shift. Spatial error analysis indicates that RF and DNN show pronounced error clustering in regions such as Michigan and Ohio. While DANN partially alleviates this issue through adversarial training, ADANN further reduces prediction errors in core agricultural zones like Illinois, underscoring its robustness under dynamic cross-regional shifts.
@@
Overall, the experiments validate the effectiveness of domain-adversarial training and entropy-guided sub-domain partitioning in TransFusion-DualDA, particularly under cross-sensor and cross-region conditions. This approach significantly reduces reliance on manual annotations by narrowing the domain gap in complex environments such as urban-rural transitions. Future work will focus on enhancing global feature extraction in high-resolution scenes to improve adaptability in dynamic remote sensing applications.
@@ 6 Challenges and Future Research Directions
Current research in agricultural intelligence faces multifaceted challenges involving data complexity, methodological adaptability, and technological limitations. Although domain adaptation (DA) techniques have shown promise in addressing cross-domain issues—such as environmental heterogeneity and annotation scarcity—they continue to face significant bottlenecks in computational efficiency, interpretability, and generalizability under complex field scenarios. This section identifies key practical bottlenecks and proposes targeted research directions.
@@ 6.1 Model Adaptability Barriers in Field Scenarios
The deployment of deep models in real-world agricultural environments is hindered by the following core issues:
1) Multi-scale spatiotemporal data fusion: Significant resolution discrepancies exist across satellite (10–100 m/pixel), UAV (0.1–1 m/pixel), and ground-level (centimeter-scale) imagery. Existing 3D convolutional approaches struggle to effectively align features across such scales.
2) Severe labeled data scarcity: Field datasets often contain fewer than 50 samples per class, particularly for rare crop diseases. Conventional transfer learning approaches often experience performance degradation exceeding 30% when training samples fall below 100.
3) Limitations in multimodal collaboration: Hyperspectral imaging in field conditions is subject to radiometric fluctuations (±30% in light intensity), inducing distributional shifts over time. Such variance compromises the effectiveness of direct multimodal fusion strategies.
@@ 6.2 Key Future Research Directions
To address these challenges, we propose four critical research directions:
(1) Lightweight Spatiotemporal Modeling Framework: Design ultra-compact network architectures with fewer than 1M parameters, incorporating physical priors (e.g., crop growth models) to guide feature extraction. Emphasize spatiotemporal decoupling to enable effective cross-scale fusion between UAV (5 cm) and satellite (10 m) imagery. Target energy consumption should be reduced to 20% of current 3D CNN baselines.
(2) Weakly Supervised Transfer Learning Paradigm: Develop pseudo-labeling mechanisms leveraging the temporal continuity of crop growth stages to generate self-supervised signals. Implement curriculum learning guided by phenological patterns to achieve over 85% classification accuracy with fewer than 100 labeled samples.
(3) Robust Multimodal Fusion System: Design spectral normalization-based feature extractors to suppress illumination-induced noise. Introduce physically informed adversarial training to ensure cross-modality consistency under lighting variation of up to ±40%. Establish sub-meter alignment protocols to address spatial inconsistencies between aerial and ground-level sensors.
(4) Privacy-Preserving Collaborative Learning: Enhance communication efficiency in federated learning frameworks to support daily updates within 10 Mbps bandwidth. Integrate differential privacy strategies compliant with GDPR, ensuring that performance degradation remains below 5% under a privacy budget of $\epsilon=2$.
To facilitate real-world deployment, we recommend the construction of a standardized open benchmark that includes over 10 representative crops across three or more climate zones. This benchmark should incorporate realistic noise factors (e.g., device error ±5%, occlusion rate 20%) to better align algorithmic development with practical robustness.
@@ 7 Conclusion
Despite recent advances, adaptive modeling in agriculture continues to face persistent challenges, particularly in addressing the complexity of spatiotemporal dynamics, annotation scarcity, and multimodal data heterogeneity. As agricultural environments become increasingly complex, the effectiveness of existing domain adaptation techniques—especially under cross-seasonal, cross-regional, and cross-modal scenarios—remains limited.
Future research should aim to develop task-specific, high-dimensional adaptive models to enhance performance in diverse agricultural applications such as crop health monitoring, yield prediction, and pest recognition. Unsupervised domain adaptation remains especially valuable in data-scarce scenarios, while multimodal fusion—including image, meteorological, and soil data—offers further potential to enhance model generalization. Moreover, source-free domain adaptation represents a critical research frontier, particularly when access to source data is restricted due to privacy or operational constraints.
Ultimately, the robustness of these models in field deployments depends on their ability to operate under unpredictable environmental conditions, across diverse crop types, and using heterogeneous sensor platforms. Overcoming these technical and operational barriers is essential to advance precision agriculture and intelligent farming. Enhancing model resilience and cross-domain adaptability will enable more accurate agricultural decision-making, yield forecasting, and disease monitoring, thus contributing to the sustainable transformation of modern agriculture.